A sample of gas has an initial pressure of 1.5 atm, an initial volume of 3.0 L, and an initial temperature of 293K. If the final pressure is 2.5 atm and the final temperature is 303K, what will the final volume be?


1
Expert's answer
2021-05-03T07:35:55-0400

P1V1/T1 = P2V2/T2

1.5 × 3.0/293 = 2.5 ×V2/303

0.0154 = 0.0083V2

= 1.86L